A 60g sample of iron reacts with 26g
of oxygen to form how many grams of iron oxide?
Express your answer using two significant figures.
1
Expert's answer
2019-06-20T04:06:08-0400

4Fe + 3O2 --> 2Fe2O3

n=m/M

n(Fe)=60/56=1.07 mol

n(O2)=26/32=0.81 mol

n(Fe)/4 : n(O2)/3 = 0.2678 : 0.2708 --> Fe is a limiting agent

n(Fe2O3)=n(Fe)/2

n(Fe2O3)=1.07/2=0.535 mol

M(Fe2O3) = 56*2 + 16*3 = 160 g/mol

m(Fe2O3)= 0.535*160 = 85.6 g


Answer: 85.6 g